A painting of mass 3.20 kg hangs on a wall. Two thin pieces of wire, each 0.250 m long, connect the painting's center to two hooks in the wall. The hooks are at the same height and are 0.330 m apart. When the painting hangs straight on the wall, how much tension is in each piece of wire? In N

The answer is not 28.97N
Answer #1

From figure,

By force balance in vertical direction,

T*sin + T*sin = W

here, T = tension in each piece of wire

W = weight of painting = m*g

W = 3.20*9.81 = 31.39 N

from figure, cos = (0.330/2)/0.250 = 0.66

then, sin = sqrt(1 - (cos)^2) = sqrt(1 - 0.66^2) = 0.75

So, T*0.75 + 0.75*T = 31.39

T = 31.39/(0.75+0.75)

T = 20.93 N

Please upvote.